heathergreen:

In order to solve for the variable in the equation 2 (4 x + 3) + 4 = 3 x minus (2 x + 4), Giovanni first applies the distributive property. Which equation is a result of this step? 8 x + 3 + 4 = 3 x minus 2 x + 4 8 x + 6 + 4 = 3 x minus 2 x minus 4 6 x + 5 + 4 = 3 x minus x + 3 6 x + 3 + 4 = 3 x minus x + 4

IamAku:

ez

IamAku:

Answer: 8x+6+4=3x-2x-4 Step-by-step explanation: the distributive property is multiplying the numbers outside of the parentheses into the inside 2×4x=8x 2×3=6 8x+6+4 Then do the same thing on the other side Which gets you 3x-2x-4 so he has the equation 8x+6+4=3x-2x-4

heathergreen:

Oh